# One-Embedding-Fits-All: Efficient Zero-Shot Time Series Forecasting by a Model Zoo

### v1

@@ Abstract
The proliferation of Time Series Foundation Models (TSFMs) has significantly advanced zero-shot forecasting, enabling predictions for unseen time series without task-specific fine-tuning. Extensive research has confirmed that no single TSFM excels universally, as different models exhibit preferences for distinct temporal patterns. This diversity suggests an opportunity: how to take advantage of the complementary abilities of TSFMs. To this end, we propose ZooCast, which characterizes each model’s distinct forecasting strengths, ZooCast can intelligently assemble current TSFMs into a model zoo that dynamically selects optimal models for different forecasting tasks. Our key innovation lies in the One-Embedding-Fits-All paradigm that constructs a unified representation space where each model in the zoo is represented by a single embedding, enabling efficient similarity matching for all tasks. Experiments demonstrate ZooCast’s strong performance on GIFT-Eval zero-shot forecasting benchmark while maintaining the efficiency of a single TSFM. In real-world scenarios with sequential model releases, the framework seamlessly adds new models for progressive accuracy gains with negligible overhead.
$\dagger$ Corresponding authors
@@ 1 Introduction
Time-series forecasting is crucial across domains, including finance omer2020financial , meteorology zahra2020transductive , industrial systems kukjin2021deep ; ming24a ; olly22multi , healthcare penfold2013use , and environmental science zahra2020transductive , enabling informed decision-making through historical pattern analysis. Recently, the emergence of Time Series Foundation Models (TSFMs) Survey-PTM has significantly advanced zero-shot forecasting, where a model predicts future values directly from a historical input window without task-specific fine-tuning. However, these independently developed TSFMs vary greatly in architecture and training methods, leading to diverse predictive strengths and weaknesses GE ; forecastpfn ; GruverFQW23 . For instance, Chronos demonstrates particular strength in high-frequency electricity data GE , while VisionTS significantly outperforms peers on cloud data full of spikes Cloud_Data . Since no single model achieves optimality universally, this naturally leads to an opportunity where the complementary strengths of different models can be combined. To this end, we introduce a model zoo, a curated collection of diverse TSFMs that leverages their complementary abilities. Such a zoo not only enhances forecasting performance beyond any single TSFM but also preserves the efficiency of zero-shot inference, providing a principled direction for building stronger forecasting systems.
How to utilize existing TSFMs in a model zoo for optimal zero-shot performance on unseen forecasting tasks? A simple solution would be enumerating all TSFMs for new tasks or naively ensembling all of them. While these approaches may achieve relatively good performance, contemporary TSFMs typically range from hundreds of megabytes to several gigabytes in size VisionTS ; TimesFM ; moment ; MOIRAI , making such exhaustive methods computationally prohibitive. We identify the key challenges for model-zoo-based forecasting: how to select among current models to enhance zero-shot performance on arbitrary forecasting tasks while maintaining efficiency, keeping computational overhead acceptable.
@@ 2 Related Work
Time Series Foundation Models. Unlike traditional deep learning approaches that focus on capturing complex temporal dependencies through sequence models  Wu2021Autoformer ; fedformer ; iTransformer ; han2024the , emerging methods leverage large language models (LLMs) for time series analysis by exploiting their heterogeneous data pretraining capabilities TimeGPT ; onefitsall ; TimeLLM ; TSFMsurvey ; TEMPO ; LLM4TS . However, the direct transferability of such approaches has been questioned due to fundamental modality gaps between time series and textual data NoLLM4TS . Recent progress in pretrained time series foundation models (TSFMs) demonstrates strong zero-shot forecasting potential. By learning generalized temporal patterns from massive multivariate corpora  forecastpfn ; metanbeats , TSFM can generate predictions through temporal context understanding with minimal adaptations moment ; TimesFM ; VisionTS ; MOIRAI ; chronos ; tabpfnv1 ; tabpfnv2 ; liu2025sundialfamilyhighlycapable . Yet these independently developed TSFMs exhibit significant architectural and training divergences, while real-world time series data vary substantially across domains, lengths, frequencies, and dimensions. Consequently, even zero-shot-capable TSFMs cannot guarantee universal superiority across all downstream tasks GE ; forecastpfn ; GruverFQW23 ; Cloud_Data , leaving practitioners without clear guidance for model selection in real-world deployments.
Model Selection Methods. Early model selection in transfer learning relied on brute-force fine-tuning, exhaustively adapting each model to the target task. While accurate, this was infeasible at scale. To reduce cost, forward-based methods emerged, approximating model suitability through lightweight inference. LEEP LEEP estimates the conditional distribution between source and target labels via a single forward pass, while LogME LogME computes a marginal likelihood to evaluate feature–label alignment for both classification and regression, though it assumes Gaussian labels and is costly in high dimensions. Despite efficiency gains, forward-based methods still depend on labeled target data and light training, making them incompatible with strict zero-shot settings. To eliminate the need for target labels or forward passes, representation-based methods compare tasks in a shared embedding space. TASK2VEC TASK2VEC pioneered this idea using Fisher information but is tied to classification architectures. TS2Vec TS2Vec introduced multi-scale temporal embeddings for time series, but its data-driven embeddings are not aligned with model behavior. SimMTM simmtm adopted masked modeling to predict missing values, which is lightweight and effective for pretraining but lacks mechanisms for transferability or model selection. TimesURL TimesURL enhanced TS2Vec with frequency-time augmentations, yet remains computationally heavy and requires post hoc model ranking. ModelSpider Spider framed model selection as a learning-to-rank task, effective in classification but requiring expensive fine-tuned labels and not extended to regression or time series forecasting. Overall, these methods remain limited by structural rigidity, task specificity, computational overhead, and lack of integrated ranking mechanisms, leaving time series zero-shot selection unresolved.
@@
Model Zoo for Machine Learning. Above challenge motivates the model zoo solution. Aggregating existing TSFMs into a zoo enables systematic modeling of inter-model differences reaugmentmodelzooguidedrl , precisely characterizing each model’s predictive specialties to guide selection for new tasks. While successfully applied in computer vision and NLP Spider ; Zoo_Graph ; CIT , adapting model zoos to time series requires novel designs to address unique challenges including temporal continuity, non-stationary distributions, multichannel correlations, and multifrequency patterns. Compared with prior approaches that either relied on costly labeled supervision or lacked mechanisms to capture transferability, recent advances move toward more efficient, forward-free selection strategies that emphasize semantic consistency, robustness to structural heterogeneity. Our work follows this direction and pioneers TSFM-based model zoos for time series, introducing a unified framework that enhances both prediction accuracy and practical usability in low-resource scenarios while maintaining cross-domain flexibility. Our designs show that the model zoo paradigm, successful in CV and NLP, also applies to time-series forecasting when appropriately adapted.
@@ 3.1 Foundation Models for Multivariate Time-Series Forecasting
Multivariate time-series forecasting aims to predict future values of multiple interdependent variables based on their historical observations. Formally, let $X=[\mathbf{x}_{1},\dots,\mathbf{x}_{T}]\in\mathbb{R}^{C\times T}$ denote the historical input with $C$ channels and length $T$, where $\mathbf{x}_{t}\in\mathbb{R}^{C}$ is the observation at time $t$. Given a forecasting horizon $H$, the task is to predict the future multivariate sequence $Y=[\mathbf{x}_{T+1},\dots,\mathbf{x}_{T+H}]\in\mathbb{R}^{C\times H}$ conditioned on $X$ and $H$, yielding predictions $\hat{Y}=[\hat{\mathbf{x}}_{T+1},\dots,\hat{\mathbf{x}}_{T+H}]$. This formulation makes explicit that the target is the entire $H$-step sequence rather than a single time point, and it reflects the need to model both temporal dependencies and cross-channel correlations.
Zero-shot forecasting represents the ability of one model to generate predictions for unseen datasets or tasks without any fine-tuning. This is supported by foundation models, large pretrained models with significant generalization capacity across tasks, e.g. TimeGPT TimeGPT , Chronos chronos , and Moirai MOIRAI . Mathematically, zero-shot forecasting can be expressed as $\hat{Y}=\phi(X)$ where $\phi$ is a pretrained foundation model from the zoo. Such models can be directly applied to new forecasting tasks, yielding universal solutions across different domains.
@@ 3.2 Zero-shot Forecasting with a Model Zoo
Despite the broad applicability of individual foundation models, their performance varies with architectures, training methodologies, and inductive biases forecastpfn ; GruverFQW23 . For instance, Chronos demonstrates particular strength in high-frequency electricity data GE , while VisionTS significantly outperforms peers on cloud data with frequent spikes Cloud_Data . This motivates the construction of a model zoo: a curated collection of diverse models, denoted as $\mathcal{Z}=\{\phi_{1},\phi_{2},\dots,\phi_{M}\}$. By taking advantage of the complementary properties of different TSFMs, model-zoo forecasting aims to improve accuracy while keeping prediction costs manageable.
@@
Single Model Selection (Selective Inference). The zoo first produces a ranking of models based on their estimated zero-shot performance for the input $X$, yielding $\mathbf{r}_{\text{final}}$. This allows the system to focus on the most promising candidates rather than evaluating all models.
Top-$K$ Ensemble Forecasting (Aggregated Prediction). Based on this ranking, predictions $\hat{Y}$ are aggregated from the top-$K$ models through averaging. This strategy leverages model diversity to achieve more robust predictions.
$\mathbf{r}_{\text{final}}=\mathrm{Rank}(X,\mathcal{Z}),\quad\hat{Y}=\tfrac{1}{K}\sum_{m\in\text{TopK}(\mathbf{r}_{\text{final}})}\phi_{m}(X,H).$ (1)
@@
$\max\;\begin{cases}\Delta P:=1-\mathcal{L}(\hat{Y})\big{/}\min\limits_{m=1,\dots,M}\mathcal{L}(\phi_{m}(X,H)),\\[6.0pt] \eta:=\mathbb{E}[\Delta P]\big{/}T_{\text{all}}.\end{cases}$ (2)
Here $\mathcal{L}(\cdot)$ denotes the forecasting loss, $\phi_{m}(X,H)$ is the $H$-step prediction of the $m$-th model, $T_{\text{all}}$ is the total runtime including both model ranking and prediction generation, and $\mathbb{E}[\cdot]$ denotes expectation over tasks. In summary, the model zoo setting enables forecasting through conditional model selection followed by top-$K$ ensemble prediction, with performance evaluated in terms of both accuracy improvement and efficiency.
3.3 Challenges with Model Zoo Forecasting
A major challenge in model-zoo-based forecasting arises from reducing the computational complexity of selection and forecasting steps while maintaining dominant performance. Contemporary TSFMs typically range from hundreds of megabytes to several gigabytes in size VisionTS ; TimesFM ; moment ; MOIRAI , making such exhaustive methods computationally prohibitive. Table 1 compares the complexity of different model selection approaches (see Appendix D.3 for a detailed description of the strategy and complexity analysis). Specifically: (1) Random simply chooses a single model, typically yielding the worst accuracy but representing the theoretical upper bound of efficiency. (2) Enumerate all evaluates every TSFM on the task, which achieves high accuracy at prohibitive cost. (3) Ensemble all averages predictions from all models, which also incurs the same heavy cost. (4) Forward-based methods attempt to reduce this cost by partial evaluation, but still require repeated forward passes per task. Although transferability methods Transfer_BaoLHZZZG19 ; LFC ; LEEP ; GBC ; OTCE ; NCE ; LogME alleviate some overhead, they remain computationally unaffordable as the zoo expands. This bottleneck mirrors challenges in related domains such as computer vision and natural language processing, and is further compounded by the expensive nature of temporal inference operations. Therefore, it is imperative to develop a lightweight and forward-free model selection mechanism.
Our solution employs a representation-based (repr-based) method, where “representation” and “embedding” are used interchangeably in this paper. The idea is to compact summaries from both models and datasets. During the precomputation phase, model embeddings are computed once. Model suitability is then estimated via efficient similarity matching between these embeddings, effectively amortizing selection costs across numerous tasks. Repr-based selection significantly reduces computational load, scales well with zoo size, and provides a generalizable measure of suitability across datasets.
@@ 4.1 Advantage Subset Characterization
Based on this finding, we construct advantage subsets that capture each model’s unique strengths. To prevent test data leakage, we build a zoo characterization set $\mathcal{D}$ containing $n$ subsequences by randomly sampling source-specific pools $\{\mathcal{D}_{i}\}$ from the pre-training datasets of each TSFM in the zoo $\{\phi_{m}\}_{m=1}^{M}$. This ensures broad representation and fairness. We first forward all models on $\mathcal{D}$ to obtain the MSE matrix $\mathbf{E}\in\mathbb{R}^{M\times n}$, where $E_{m,i}$ is the MSE of $\phi_{m}$ on $\mathbf{x}_{i}$. Let the inter-model error variance on sample $\mathbf{x}_{i}$ be $\sigma_{i}\;=\;\operatorname{std}\!\big{(}\{E_{m,i}\}_{m=1}^{M}\big{)}$, and let $\bar{\sigma}$ and $\hat{\sigma}$ be the mean and standard deviation of $\{\sigma_{i}\}$. We define a per-sample advantage score and the advantage subset for model $\phi_{m}$ as
$s_{m,i}\;=\;\left(\frac{1}{M-1}\sum_{k\neq m}E_{k,i}\;-\;E_{m,i}\right)\cdot\frac{\sigma_{i}-\bar{\sigma}}{\hat{\sigma}},\quad\mathcal{D}_{m}^{\text{adv}}\;=\;\left\{\,\mathbf{x}_{i}\in\mathcal{D}\;\middle|\;s_{m,i}\;>\;\tau\,\right\},$ (3)
where $\tau$ is an adaptive threshold. This criterion jointly encodes model-specific advantage and sample discriminability via $\sigma_{i}$. The precomputed matrix $\mathbf{E}$ enables efficient updates: adding a new model only appends one row to $\mathbf{E}$ and recomputes its $\{s_{m,i}\}$ without touching existing entries. The cardinality $d_{m}\;=\;\big{|}\mathcal{D}_{m}^{\text{adv}}\big{|}$ is used to derive the model weight $w_{m}$ in selection (details in Appendix D.1).

### v2

@@ Abstract
The proliferation of Time Series Foundation Models (TSFMs) has significantly advanced zero-shot forecasting, enabling predictions for unseen time series without task-specific fine-tuning. Extensive research has confirmed that no single TSFM excels universally, as different models exhibit preferences for distinct temporal patterns. This diversity suggests an opportunity: how to take advantage of the complementary abilities of TSFMs. To this end, we propose ZooCast, which characterizes each model’s distinct forecasting strengths. ZooCast can intelligently assemble current TSFMs into a model zoo that dynamically selects optimal models for different forecasting tasks. Our key innovation lies in the One-Embedding-Fits-All paradigm that constructs a unified representation space where each model in the zoo is represented by a single embedding, enabling efficient similarity matching for all tasks. Experiments demonstrate ZooCast’s strong performance on the GIFT-Eval zero-shot forecasting benchmark while maintaining the efficiency of a single TSFM. In real-world scenarios with sequential model releases, the framework seamlessly adds new models for progressive accuracy gains with negligible overhead.
@@ 1 Introduction
Time-series forecasting is crucial across domains, including finance (Sezer et al., 2020), meteorology (Karevan & Suykens, 2020), industrial systems (Choi et al., 2021; Jin et al., 2024a; Styles et al., 2022), healthcare (Penfold & Zhang, 2013), and environmental science (Karevan & Suykens, 2020), enabling informed decision-making through historical pattern analysis. Recently, the emergence of Time Series Foundation Models (TSFMs) (Ma et al., 2024) has significantly advanced zero-shot forecasting, where a model predicts future values directly from a historical input window without task-specific fine-tuning. However, these independently developed TSFMs vary greatly in architecture and training methods, leading to diverse predictive strengths and weaknesses (Aksu et al., 2024; Dooley et al., 2023; Gruver et al., 2023). For instance, Chronos demonstrates particular strength in high-frequency electricity data (Aksu et al., 2024), while VisionTS significantly outperforms peers on cloud data full of spikes (Toner et al., 2025). Since no single model achieves optimality universally, this naturally leads to an opportunity where the complementary strengths of different models can be combined. To this end, we introduce a model zoo, a curated collection of diverse TSFMs that leverages their complementary abilities. Such a zoo not only enhances forecasting performance beyond any single TSFM but also preserves the efficiency of zero-shot inference, providing a principled direction for building stronger forecasting systems.
How to utilize existing TSFMs in a model zoo for optimal zero-shot performance on unseen forecasting tasks? A simple solution would be enumerating all TSFMs for new tasks or naively ensembling all of them. While these approaches may achieve relatively good performance, contemporary TSFMs typically range from hundreds of megabytes to several gigabytes in size (Chen et al., 2025; Das et al., 2024; Goswami et al., 2024; Woo et al., 2024), making such exhaustive methods computationally prohibitive. We identify the key challenges for model-zoo-based forecasting: how to select among current models to enhance zero-shot performance on arbitrary forecasting tasks while maintaining efficiency, keeping computational overhead acceptable.
@@ 2 Related Work
Time Series Foundation Models. Traditional deep learning approaches model temporal dependencies through sequence architectures (Wu et al., 2021; Zhou et al., 2022; Liu et al., 2024; Han et al., 2024), while recent methods adapt LLMs for time series by leveraging heterogeneous pretraining (Garza & Canseco, 2023; Zhou et al., 2023; Jin et al., 2024b; Liang et al., 2024; Cao et al., 2024; Bian et al., 2024). Yet direct transfer remains limited by modality gaps (Tan et al., 2024). Pretrained TSFMs (Dooley et al., 2023; Oreshkin et al., 2021) have shown strong zero-shot forecasting ability with minimal adaptation (Goswami et al., 2024; Das et al., 2024; Chen et al., 2025; Woo et al., 2024; Ansari et al., 2024; Hoo et al., 2024; 2025; Liu et al., 2025), but they differ substantially in architecture and training, while time series tasks vary widely across domains and scales. Thus, even advanced TSFMs cannot ensure universal superiority (Aksu et al., 2024; Dooley et al., 2023; Gruver et al., 2023; Toner et al., 2025), leaving model selection unresolved in real-world deployments.
Model Selection Methods. Early transfer learning relied on brute-force fine-tuning, accurate but infeasible at scale. Forward-based methods such as LEEP (Nguyen et al., 2020) and LogME (You et al., 2021) improved efficiency but still required labels and light training, limiting strict zero-shot use. Representation-based approaches, e.g., TASK2VEC (Achille et al., 2019), TS2Vec (Yue et al., 2022), SimMTM (Dong et al., 2023), and TimesURL (Liu & Chen, 2024), map tasks into embeddings but either depend on classification settings, lack transferability guarantees, or remain costly. Learning-to-rank solutions like ModelSpider (Zhang et al., 2023) achieve accuracy with labels but are not extended to regression or forecasting. Overall, current methods suffer from rigidity, task specificity, computational overhead, and missing integrated ranking, leaving time series zero-shot selection an open problem. (See Appendix C for detailed discussion.)
@@
Model Zoo for Machine Learning. Above challenge motivates the model zoo solution. Model zoos aggregate diverse models to capture inter-model differences (Yuan et al., 2025), enabling task-specific guidance. They have succeeded in CV and NLP (Zhang et al., 2023; Li et al., 2024; Zhang et al., 2025b), but time series require designs addressing temporal continuity, non-stationarity, multichannel and multifrequency structures. Compared with prior approaches that either relied on costly labeled supervision or lacked mechanisms to capture transferability, recent advances move toward more efficient, forward-free selection strategies that emphasize semantic consistency, robustness to structural heterogeneity. Our work follows this direction and pioneers TSFM-based model zoos for time series, introducing a unified framework that enhances both prediction accuracy and practical usability in low-resource scenarios while maintaining cross-domain flexibility.
@@ 3.1 Foundation Models for Multivariate Time-Series Forecasting
Multivariate time-series forecasting aims to predict future values of multiple interdependent variables based on their historical observations. Formally, let ${\bm{X}}=[{\bm{x}}_{1},\dots,{\bm{x}}_{T}]\in\mathbb{R}^{C\times T}$ denote the historical input with $C$ channels and length $T$, where ${\bm{x}}_{t}\in\mathbb{R}^{C}$ is the observation at time $t$. Given a forecasting horizon $H$, the task is to predict the future multivariate sequence ${\bm{Y}}=[{\bm{x}}_{T+1},\dots,{\bm{x}}_{T+H}]\in\mathbb{R}^{C\times H}$ conditioned on $\bm{X}$} and $H$, yielding predictions $\hat{{\bm{Y}}}=[\hat{{\bm{x}}}_{T+1},\dots,\hat{{\bm{x}}}_{T+H}]$. This formulation makes explicit that the target is the entire $H$-step sequence rather than a single time point, and it reflects the need to model both temporal dependencies and cross-channel correlations.
Zero-shot forecasting represents the ability of one model to generate predictions for unseen datasets or tasks without any fine-tuning. This is supported by foundation models, large pretrained models with significant generalization capacity across tasks, e.g. TimeGPT (Garza & Canseco, 2023), Chronos (Ansari et al., 2024), and Moirai (Woo et al., 2024). Mathematically, zero-shot forecasting can be expressed as $\hat{{\bm{Y}}}=\phi({\bm{X}},H)$ where $\phi$ is a pretrained foundation model. Such models can be directly applied to new forecasting tasks, yielding universal solutions across different domains.
@@ 3.2 Zero-shot Forecasting with a Model Zoo
Despite the broad applicability of individual foundation models, their performance varies with architectures, training methodologies, and inductive biases (Dooley et al., 2023; Gruver et al., 2023). For instance, Chronos demonstrates particular strength in high-frequency electricity data (Aksu et al., 2024), while VisionTS significantly outperforms peers on cloud data with frequent spikes (Toner et al., 2025). This motivates the construction of a model zoo: a curated collection of diverse models, denoted as $\mathcal{Z}=\{\phi_{1},\phi_{2},\dots,\phi_{M}\}$. By taking advantage of the complementary properties of different TSFMs, model-zoo forecasting aims to improve accuracy while keeping prediction costs manageable.
@@
Single Model Selection (Selective Inference). The zoo first produces a ranking of models based on their estimated zero-shot performance for the input $\bm{X}$}, yielding ${\bm{r}}_{\text{final}}$. This allows the system to focus on the most promising candidates rather than evaluating all models.
Top-$K$ Ensemble Forecasting (Aggregated Prediction). Based on this ranking, predictions $\hat{{\bm{Y}}}$ are aggregated from the top-$K$ models through averaging. This strategy leverages model diversity to achieve more robust predictions.
${\bm{r}}_{\text{final}}=\mathrm{Rank}({\bm{X}},\mathcal{Z}),\quad\hat{{\bm{Y}}}=\tfrac{1}{K}\sum_{m\in\text{TopK}({\bm{r}}_{\text{final}})}\phi_{m}({\bm{X}},H).$ (1)
@@
$\max\;\begin{cases}\Delta P:=1-\mathcal{L}(\hat{{\bm{Y}}})\big/\min\limits_{m=1,\dots,M}\mathcal{L}(\phi_{m}({\bm{X}},H)),\\[6.0pt] \eta:=\mathbb{E}[\Delta P]\big/T_{\text{all}}.\end{cases}$ (2)
Here $\mathcal{L}(\cdot)$ denotes the forecasting loss, $\phi_{m}({\bm{X}},H)$ is the $H$-step prediction of the $m$-th model, $T_{\text{all}}$ is the total runtime including both model ranking and prediction generation, and $\mathbb{E}[\cdot]$ denotes expectation over tasks. In summary, the model zoo setting enables forecasting through conditional model selection followed by top-$K$ ensemble prediction, with performance evaluated in terms of both accuracy improvement and efficiency.
@@
3.3 Challenges with Model Zoo Forecasting
A key challenge in model-zoo-based forecasting is balancing efficiency and accuracy. Contemporary TSFMs often occupy hundreds of megabytes to gigabytes (Chen et al., 2025; Das et al., 2024; Goswami et al., 2024; Woo et al., 2024), making exhaustive evaluation computationally prohibitive. Table 1 summarizes the complexity of different strategies (details in Appendix F.3). (1) Random selection offers maximal efficiency but worst accuracy. (2) Enumerate all evaluates every TSFM for high accuracy at prohibitive cost. (3) Ensemble all averages predictions with similar expense. (4) Forward-based methods reduce cost through partial evaluation but still require repeated passes. Although transferability methods (Bao et al., 2019; Deshpande et al., 2021; Nguyen et al., 2020; Pándy et al., 2022; Tan et al., 2021; Tran et al., 2019; You et al., 2021) alleviate some overhead, they remain computationally unaffordable as the zoo expands. This bottleneck mirrors challenges in related domains such as computer vision and natural language processing, and is further compounded by the expensive nature of temporal inference operations. Therefore, it is imperative to develop a lightweight and forward-free model selection mechanism.
Our solution adopts a representation-based (repr-based) approach, using “representation” and “embedding” interchangeably. Models and datasets are summarized into embeddings, with model embeddings precomputed once. Suitability is then estimated via efficient similarity matching, amortizing selection cost across tasks. Repr-based selection cuts computational load, scales with zoo size, and offers a transferable suitability measure across datasets.
@@ 4.1 Advantage Subset Characterization
Based on this finding, we construct advantage subsets that capture each model’s unique strengths. To prevent test data leakage, we build a zoo characterization set $\mathcal{D}$ containing $n$ subsequences by randomly sampling source-specific pools $\{\mathcal{D}_{i}\}$ from the pre-training datasets of each TSFM in the zoo $\{\phi_{m}\}_{m=1}^{M}$. The advantage subset for model $\phi_{m}$ is defined as:
$\mathcal{D}_{m}^{\text{adv}}\;=\;\left\{\,{\bm{x}}_{i}\in\mathcal{D}\;\middle|\;s_{m,i}\;>\;\tau\,\right\}.$ (3)
Here, $s_{m,i}$ denotes the advantage score and $\tau$ is an adaptive threshold. The detailed computation and interpretation of the advantage score are provided in Appendix E.1. This design allows each model’s unique strengths to be efficiently profiled while maintaining scalability.
4.2 Model-task Representations Co-Embedding
Even with model-specific subsets, another challenge is embedding tasks and models into a shared space despite non-stationarity and multi-channel structure. We address this with a co-embedding extractor that unifies them for efficient similarity-based matching.